OpenStudy (anonymous):

A student wearing frictionless in-line skates on a horizontal surface is pushed by a friend with a constant force of 45 N. How far must the student be pushed, starting from rest, so that her final kinetic energy is 352 J?

OpenStudy (ujjwal):

we have F=ma, i.e. a=F/m according to info given in question a=45/m.....(1) again, (1/2)mv^2 = 352 or, v^2= 352 * 2/m.....(2) Also we have v^2=2as.... (3) (starting from rest) solving (1), (2) and (3) we get s=7.82 m hence the student must be pushed 7.82 m..

OpenStudy (anonymous):

What does "a" stand for?

OpenStudy (ujjwal):

acceleration

OpenStudy (anonymous):

alright Thank You! :D

OpenStudy (ujjwal):

you are most welcome.. :)

OpenStudy (anonymous):

Another method: By using principal of conservation of energy: workdone=change in kinetic energy F*d=(K.E)final-0; d=(K.E)final/F=352/45=7.82m

OpenStudy (anonymous):

Thank You!